OpenStudy (megannicole51):

Does the series converge or diverge? sqrt(n^4+1)/5n^2 This one i actually need help on...what test do I use for this one? @LolWolf

OpenStudy (anonymous):

Try using a direct comparison. Here: \[ \frac{\sqrt{n^4+1}}{5n^2}=\frac{1}{5}\sqrt{\frac{n^4}{n^4}+\frac{1}{n^4}}>\;? \]

OpenStudy (megannicole51):

p>1 converges

OpenStudy (megannicole51):

i dont see how you got the fractions under the sqrt though

OpenStudy (megannicole51):

jk i see it now

OpenStudy (megannicole51):

sorry its been a long night!

OpenStudy (anonymous):

No, it doesn't quite converge. Note that:\[ \sqrt{\frac{n^4}{n^4}+\frac{1}{n^4}}=\sqrt{1+\frac{1}{n^4}}>1 \] And, clearly: \[ 1+1+1+1+1+1+\cdots=\infty \]

OpenStudy (megannicole51):

but since its larger than 1, because of the p-series it converges? I'm sorry im not quite seeing it

OpenStudy (anonymous):

Remember that the p-series refers to the exponent of the inverse: \[ \sum_n\frac{1}{n^p} \]This is not an exponent, this is the actual number: \[ \sum_n\sqrt{1+\frac{1}{n^4}}>\sum_n 1=\infty \]Sum 1 infinitely many times and you should have infinity.

OpenStudy (megannicole51):

oooh okay! i get it now...thank you:)